### v1

@@ 5.1.1 Gate Count
In the NISQ era, quantum gate noise is non-negligible. To minimize overall circuit noise, reducing the number of quantum gates in the mapped circuit is essential. It should be noted that various studies [107, 191, 144, 130, 217, 38, 216, 172, 113, 109, 213, 212, 77, 148, 114, 145] may differ slightly in their approach to optimizing quantum gate count. These approaches can be broadly categorized into optimizing the total quantum gate count, the number of added quantum gates, SWAP gate count, and CNOT gate count. Notably, quantum gates added in the compiling stage are mainly the SWAP gates, and SWAP gates require only CNOT gates to implement. Thus, the above-mentioned objectives are unified as optimizing gate count.
5.1.2 Execution Time
Circuit execution time represents time performance, reflecting circuit execution efficiency, and also affects circuit execution results. Due to the physical qubit decoherence problem, longer circuit runtime increases the likelihood of qubit decoherence, leading to errors in circuit execution results. Consequently, several studies [135, 144, 206] consider circuit execution time during circuit mapping, aiming to minimize mapped circuit execution time while satisfying connectivity constraints.
5.1.3 Circuit Depth
@@
This optimization objective usually appears in conjunction with the number of quantum gates. It reflects both the number of quantum gates and, to some extent, circuit execution time. Generally, deeper circuits imply longer execution time and increased likelihood of decoherence. Consequently, circuit depth is often included in the optimization objectives of quantum circuit mapping [107, 217, 38, 138, 113, 171, 155, 114].
5.1.4 Fidelity
@@
Fidelity measures the similarity between two quantum states and can be used to assess the accuracy of circuit execution results by comparing the output quantum state to the theoretically expected state. However, direct qubit state measurement is challenging, making fidelity calculation on quantum computers difficult. Therefore, most studies [178, 135, 144, 113, 114, 145] utilize the PST (Probability of a Successful Trial) index to reflect fidelity, representing the probability of obtaining correct execution results, which can be directly calculated from measurement outcomes. Alternatively, [13] employs Kraus operators to simulate quantum gate noise models and qubit decoherence time, enabling direct calculation and optimization of execution result fidelity.
5.1.5 Turnaround Time
@@
Nowadays, companies like IBM have released their quantum resources to the public. People can use their quantum computers for free on their quantum cloud. However, with the growing need for quantum resources, one may queue up on the cloud for a long time to run their quantum circuits. To fully utilize the limited quantum resources, Das et al.[40] propose multi-programming, i.e. executing multiple quantum circuits on a quantum computer concurrently. Wu et al.[194] further formulate it as a scheduling problem. An important objective in this problem is to minimize the average turnaround time (i.e. time cost from submission to completion of a quantum job), given the coupling graph, noise conditions, and quantum jobs in the queue.
5.2 Quubit Mapping and Routing Methods
@@ 5.2.1 Optimal Solution
Siraichi et al.[172] present a method to obtain an exact solution based on dynamic programming. Assuming a circuit with $n$ qubits and $m$ two-qubit gates, the solution to the subproblem is defined as $S(l,i)$, representing the minimum number of SWAP gates required to satisfy the constraints of the first $i$ two-qubit gates with the final qubit mapping being $l$. The solution to the original problem can thus be derived from various subproblem solutions. While this method yields the optimal solution, its time complexity of $\mathcal{O}(n!^{2}\cdot n\cdot m)$ significantly limits its scalability.
5.2.2 Heuristic Algorithms
@@
This approach, prevalent in existing work, generally plans SWAP gate paths using noise conditions and inter-qubit distances to find near-optimal solutions efficiently. As quantum circuits can be converted into DAGs, some studies apply classical graph algorithms to solve this problem. Zulehner et al.[217] calculate local optimal solutions between circuit layers using the A* algorithm, incorporating a look-ahead mechanism to consider subsequent two-qubit gate impacts, approximating the global optimal solution. Zhang et al.[206] also employ the A* algorithm but consider individual gate execution time to optimize overall circuit execution time. Tannu et al.[178] utilize the Dijkstra algorithm for an approximate solution. To further reduce search space and improve solution quality, Li et al.[107] introduce a scoring function guiding heuristic search, determined by the weighted sum of distances between all two-qubit gates in current and look-ahead layers. The heuristic search selects SWAP gates, minimizing this scoring function. Building on this, several studies[216, 144, 113, 148, 145] have proposed improvements. Zhu et al.[216] first select SWAP gates reducing current layer qubit distances as candidates, then choose those maximally reducing look-ahead layer qubit distances. Niu et al.[144] incorporate noise and execution time terms into the distance function and introduce BRIDGE gates as SWAP alternatives. In[145], they further consider SWAP and BRIDGE gate noise, enhancing scoring function realism. Liu et al.[113] incorporate circuit optimization, reducing CNOT gate count through two-qubit block optimization and multiple two-qubit gate cancellation. Park et al.[148] add distance-based weights to each two-qubit gate term in the scoring function, assigning lower weights to longer distances.
5.2.3 Reduction
@@
The quantum circuit mapping problem can be reduced to classical problems and solved using the corresponding solvers. Murali et al.[135], Wille et al.[191], and Molavi et al.[130] transform this problem into a Boolean Satisfiability Problem, while respectively employing SMT (Satisfiability Modulo Theories) solver, SAT solver, and MaxSAT solver to obtain solutions. Additionally, this problem can be reduced to a binary integer programming problem and solved using the commercial solver CPLEX[138]. As quantum circuits can be converted into DAGs, the initial mapping problem can be viewed as a Subgraph Isomorphism Problem. Ash-Saki et al.[13] first determine the graph structure of the initial mapping, then select the least noisy subgraph from all isomorphic subgraphs on the quantum computer topology as the initial mapping. Li et al.[109] consider the subgraph formed by all two-qubit gates in the top layer and find an isomorphic subgraph as the initial mapping. Due to the difficulty of achieving accurate subgraph matching, Park et al.[148] only require finding an approximation of the target subgraph structure as the initial mapping, while minimizing the difference between the two.
5.2.4 Machine Learning
@@
The quantum circuit mapping problem can be viewed as a decision problem, making it amenable to reinforcement learning approaches. Pozzi et al.[155] define the state as the current circuit layout and target mapping, actions as a series of parallel executable SWAP gates, and rewards determined by the number of plannable gates after action. They employ a simulated annealing algorithm for action search, with the action value function approximated by the Deep Q Learning algorithm. Sinha et al.[171] augment the state with unplanned gates and utilize the Monte Carlo tree search algorithm for action search, estimating the value function with a graph neural network. Reinforcement learning can also be applied exclusively to initial mapping. Huang et al.[77] employ a multi-head attention mechanism to encode the circuit and decode it to obtain the initial mapping, then recursively extract sub-circuit structures and use the A* algorithm for routing. They use the number of gates added after circuit mapping as the reward.
6 Summary and Outlook
@@
This survey comprehensively summarizes methods related to quantum logic circuit design and compilation optimization, encompassing quantum circuit representation, synthesis, optimization, qubit mapping, and routing. To execute quantum programs on NISQ-era quantum processors, quantum algorithms must address various constraints such as limited built-in gate sets, qubit coherence time, noise conditions, etc. Consequently, efficient transformation of quantum algorithms into executable quantum programs remains a core challenge in demonstrating quantum supremacy. The past decade, particularly with advancements in artificial intelligence, has witnessed significant progress in quantum logic circuit design and compilation optimization. Existing AI algorithms address each sub-step of the aforementioned processes separately. However, when the entire process is divided into multiple sub-modules processed by different algorithms, achieving unified goals between modules often results in suboptimal overall performance.
To maximize the potential and value of NISQ-era quantum processors and explore truly meaningful applications with quantum supremacy (distinct from random sampling tasks specifically designed for quantum computers [12, 197, 211]), more intelligent, efficient, and accurate logic circuit design and compilation optimization methods are required. Specifically, we propose that relevant methods should meet three main criteria: First, synthesize optimal logic circuits corresponding to quantum algorithms based on the quantum processors’ built-in gate set. Given the current lack of a unified execution standard for quantum processors, algorithms must adapt to diverse built-in gate sets when synthesizing quantum logic circuits. Second, conduct qubit and coupler selection, mapping, and routing based on quantum processor calibration data. Algorithms should make real-time decisions based on dynamic qubit quality. Finally, balance theoretical accuracy and noise. Achieve optimal execution results on quantum processors while judiciously sacrificing theoretical accuracy. For instance, reducing circuit depth may decrease theoretical accuracy but could lead to better practical results by reducing circuit noise and minimizing the error between actual results and theoretical accuracy during execution.
We observe that optimization goals of individual sub-modules may not always align with the overall objective. Therefore, exploring end-to-end circuit synthesis optimization algorithms from quantum algorithms to executable quantum programs is a key future direction. Leveraging artificial intelligence’s powerful perception and decision-making capabilities, algorithms can consistently optimize for the overall goal of obtaining optimal practical results on quantum processors throughout the entire logic circuit design and compilation optimization process.
Currently, quantum computing resources are primarily deployed as cloud-based quantum platforms [82, 79]. Designing and deploying an end-to-end algorithm on a quantum cloud platform would greatly facilitate the exploration and development of practically valuable quantum algorithms. Users would only need to submit their designed quantum algorithms, and the artificial intelligence algorithm could obtain optimal experimental results based on real-time calibrated data. This approach would enhance algorithm performance and efficiency, reduce additional human costs following quantum algorithm design, and provide a new paradigm for quantum computing implementation. We anticipate increased engagement from the quantum computing community, particularly in quantum logic circuit design and compilation optimization, to collectively explore the vast potential of quantum computing.
[…]
This work was supported by the Key Project of the National Natural Science Foundation of China (Grant No: 92370201), Excellent Youth Fund Project of the National Natural Science Foundation of China (Grant No: 62222607).

### v2

@@ V-A1 Gate Count
In the NISQ era, quantum gate noise is non-negligible. To minimize overall circuit noise, reducing the number of quantum gates in the mapped circuit is essential. It should be noted that various studies [59, 43, 37, 44, 42, 217, 218, 41, 38, 219, 220, 221, 47, 222, 39, 40] may differ slightly in their approach to optimizing quantum gate count. These approaches can be broadly categorized into optimizing the total gate count, the number of added gates, SWAP gate count, and CNOT gate count. Notably, quantum gates added in the compiling stage are mainly the SWAP gates, and SWAP gates require only CNOT gates to implement. Thus, the objectives mentioned above are unified as optimizing gate count.
V-A2 Circuit Depth
@@
This optimization objective reflects both the number of quantum gates and, to some extent, circuit execution time. Generally, deeper circuits imply longer execution time and increased likelihood of decoherence. Consequently, circuit depth is often included in the optimization objectives of quantum circuit mapping [59, 42, 217, 223, 38, 46, 45, 39].
V-A3 Execution Time
A critical metric that correlates with both gate count and circuit depth. This is particularly important due to qubit decoherence, as a longer circuit runtime increases the probability of quantum states degrading, which leads to execution errors. Consequently, several studies [36, 37, 224] minimize the total execution time of the mapped circuit while satisfying the hardware’s connectivity constraints.
V-A4 Fidelity
@@
A measure of similarity between quantum states is used to quantify the accuracy of circuit execution results by comparing the output state with the theoretically expected one. However, calculating fidelity directly on a quantum computer is challenging due to the complexity of measuring a complete qubit state. For this reason, many studies employ the Probability of a Successful Trial (PST) to reflect fidelity, as it can be directly computed from measurement outcomes [35, 36, 37, 38, 39, 40]. Alternatively, [225] employed Kraus operators to model quantum gate noise and qubit decoherence, enabling direct calculation and optimization of the execution fidelity through classical simulation.
V-A5 Turnaround Time
@@
The increasing demand for quantum resources has led to significant queueing delays on cloud platforms like those offered by IBM. To address this, multi-programming [226] has been proposed as a way to execute multiple quantum circuits concurrently on a single quantum processor, thereby improving the utilization of limited quantum resources. This approach is further formulated as a scheduling problem [227]. An important objective in this problem is to minimize the average turnaround time—the total time from the submission to the completion of a quantum job—given the quantum computer’s coupling graph, noise conditions, and the queue of submitted jobs.
V-B Qubit Mapping and Routing Methods
@@ V-B1 Optimal Solution
[41] presented a method to obtain an exact solution based on dynamic programming. Assuming a circuit with $n$ qubits and $m$ two-qubit gates, the solution to the subproblem is defined as $S(l,i)$, representing the minimum number of SWAP gates required to satisfy the constraints of the first $i$ two-qubit gates with the final qubit mapping being $l$. The solution to the original problem can thus be derived from various subproblem solutions. While this method yields the optimal solution, its time complexity of $\mathcal{O}(n!^{2}\cdot n\cdot m)$ significantly limits its scalability.
V-B2 Heuristic Algorithms
@@
It generally plans SWAP gate paths using noise conditions and inter-qubit distances to find near-optimal solutions efficiently. As quantum circuits can be converted into DAGs, some studies apply classical graph algorithms to solve this problem. [42] calculated local optimal solutions between circuit layers using the A algorithm, incorporating a look-ahead mechanism to consider subsequent two-qubit gate impacts, approximating the global optimal solution. The authors in [224] also employed the A algorithm but consider individual gate execution time to optimize overall circuit execution time. [35] utilized the Dijkstra algorithm for an approximate solution. To further reduce search space and improve solution quality, [59] introduced a scoring function guiding heuristic search, determined by the weighted sum of distances between all two-qubit gates in current and look-ahead layers. The heuristic search selects SWAP gates, minimizing this scoring function. Building on this, several studies [218, 37, 38, 222, 40] have proposed improvements. [218] first selected SWAP gates that reduce current layer qubit distances as candidates, then chose those maximally reducing look-ahead layer qubit distances. [37] incorporated noise and execution time terms into the distance function and introduced BRIDGE gates as SWAP alternatives. In [40], they further considered SWAP and BRIDGE gate noise, enhancing scoring function realism. [38] incorporated circuit optimization, reducing CNOT gate count through two-qubit block optimization and multiple two-qubit gate cancellation. [222] added distance-based weights to each two-qubit gate term in the scoring function, assigning lower weights to longer distances.
V-B3 Reduction
@@
The quantum circuit mapping problem can be reduced to classical problems and solved using the corresponding solvers. A series of works [36, 43, 44] transformed it into a Boolean Satisfiability Problem, while respectively employing SMT (Satisfiability Modulo Theories) solver, SAT solver, and MaxSAT solver to obtain solutions. Additionally, this problem can be reduced to a binary integer programming problem and solved using the commercial solver CPLEX [223]. As quantum circuits can be converted into DAGs, the initial mapping problem can be viewed as a Subgraph Isomorphism Problem. [225] first determined the graph structure of the initial mapping, then selected the least noisy subgraph from all isomorphic subgraphs on the quantum computer topology as the initial mapping. [219] considered the subgraph formed by all two-qubit gates in the top layer and found an isomorphic subgraph as the initial mapping. Due to the difficulty of achieving accurate subgraph matching, [222] only required finding an approximation of the target subgraph structure as the initial mapping, while minimizing the difference between the two.
V-B4 Machine Learning
@@
The quantum circuit mapping problem can be viewed as a decision problem, making it amenable to reinforcement learning approaches. [45] defined the state as the current circuit layout and target mapping, actions as a series of parallel executable SWAP gates, and rewards determined by the number of plannable gates after action. They employ a simulated annealing algorithm for action search, with the action value function approximated by the Deep Q Learning algorithm. [46] augmented the state with unplanned gates and utilized the Monte Carlo tree search algorithm for action search, estimating the value function with a graph neural network. Reinforcement learning can also be applied exclusively to initial mapping. [47] employed a multi-head attention mechanism to encode the circuit and decode it to obtain the initial mapping, then recursively extract sub-circuit structures and use the A algorithm for routing. They use the number of gates added after circuit mapping as the reward.
VI Conclusion and Outlook
@@
This survey reviews methods related to quantum logic circuit design and compilation optimization, covering quantum circuit representation, synthesis, qubit mapping, and routing. To execute quantum programs on NISQ-era processors, algorithms must address constraints, e.g., limited gate sets, qubit coherence times, and noise. Efficiently transforming quantum algorithms into executable programs remains a critical challenge for demonstrating quantum supremacy. Over the past decade, advancements in artificial intelligence have significantly impacted quantum logic circuit design and optimization, yet existing AI algorithms address each step independently. When the process is divided into sub-modules handled by separate algorithms, achieving unified goals often leads to suboptimal performance. Therefore, future research should focus on end-to-end circuit synthesis optimization—from quantum algorithms to executable programs—by leveraging AI’s ability to optimize the entire process holistically.
Quantum computing resources are currently accessible primarily through cloud-based platforms [228, 229], with future deployment anticipated to shift towards distributed quantum computers [230]. Developing end-to-end algorithms on these platforms will enable seamless integration of quantum logic circuit design with real-time calibration data, facilitating the optimization of practical quantum algorithms. It will enhance quantum algorithm performance, reduce human intervention, and provide a novel paradigm for quantum computing implementation. As the quantum computing community advances towards FTQC, the scope of these optimization methods will extend beyond NISQ-era constraints, enabling the realization of fault-tolerant quantum algorithms with significantly improved scalability and robustness.